OpenStudy (anonymous):

bob's money is 2 1/3 times becky's. bob's money is also 2 3/5 times charlie's. what fraction of becky's money is charlie's?

OpenStudy (anonymous):

half of beckys,my medla plz

OpenStudy (anonymous):

equation!!

OpenStudy (anonymous):

x is bob cash,b is beckys money and c is charlis,so x= 2*1/3b=2*2/3c,then remove x and solve for c

OpenStudy (anonymous):

my medal plz

OpenStudy (anonymous):

Set up a system of equations where Bob is A, Becky is B, and Charlie is C: A = (7/3)B A = (13/5)C Rearrange for C and substitute to solve for C: C = (5/13)A C= (5/13)(7/3)B C = (35/39)B

OpenStudy (anonymous):

Mcmadlala, you set up your equation with "2 1/3" as "2*1/3." It should be 2 plus 1/3, not 2 times 1/3. Also, you set your second fraction to be "2 2/3" when it should really be "2 3/5." I know that you did not just make typos, either, because the way that you wrote the equation gives the answer of one half of Becky's, but I believe that the answer is actually 35/39 of Becky's.